# GAGS: Granularity-Aware Feature Distillation for Language Gaussian Splatting

### v1

@@ 1 Introduction
A noticeable challenge of learning such open-vocabulary features for 3D fields lies in the multiview inconsistency of the extracted 2D features. As shown in  Fig. 2, given a specific 3D location, it can be described as “desktop”, “bowl”, and “Ramen” at the same time according to the different granularities. In this case, the extracted 2D features show no multiview consistency on different viewpoints. Brute-force training a 3D feature field from inconsistent 2D features leads to averaged and degenerated 3D features. LangSplat [30] resolves this problem by learning three different feature fields corresponding to different segmentation granularity of SAM [20]. However, learning three different fields increases both training memory and time cost and requires the query texts to be compared with three different features. N2F2 [2] proposes to merge the features of different granularity by selecting the most activated features within a set of manually predefined descriptions. In this case, the performance is strongly related to the quality of these predefined descriptions. If the granularity of query descriptions differs from these predefined descriptions, the performance degenerates severely.
Figure 3: In comparison to SAM with default prompt points, we associate the prompt point density with distance to the object, which greatly enhances the consistency of multi-view segmentation.
@@
First, we improve the segmentation granularity to improve the consistency. In the CLIP feature extraction, we need to apply the SAM to segment images into a region and extract a CLIP feature for the zoomed segmented region. SAM may have multiple possible solutions to segment an image, which introduces inconsistency for segmentation from different viewpoints. As shown in  Fig. 3, we improve the consistency of SAM segmentation results among multiview images by introducing a granularity-aware prompting strategy. SAM relies on the prompt points to determine the segmentation granularity. To keep the multiview consistency, images with far-away viewpoints should have dense prompt points while images with nearby viewpoints should have sparse prompt points. Thus, we associate the prompt point density with the camera distance to the subject by utilizing the reconstructed 3D GS. We found that this adaptive sampling effectively results in consistent segmentation results on multiview images and improves the multiview consistency of CLIP features.
@@ 2.2 SAM
As one of the most impressive foundational vision models, Meta’s SAM [20] has demonstrated exceptional zero-shot 2D segmentation capabilities. SAM supports flexible prompts, allowing it to generate multi-level masks for target objects based on inputs such as points, bounding boxes, masks, and text. Additionally, SAM offers the capability to automatically segment the entire image using predefined dense grid point prompts. Numerous methods have since emerged to extend SAM’s capabilities to 3D space. Anything-3D [34] elevates SAM’s segmentation to 3D, while Part123 [24] uses SAM’s segmentation masks to reconstruct 3D models with high-quality segmented parts. Feature 3DGS [46] and Gaussian Grouping [42] integrate SAM’s features and object segmentation results into 3D Gaussians, achieving high-quality segmentation of novel views and 3D scenes. Semantic Gaussian [13] utilizes various prompts to obtain instance-level segmentations for feature extraction. We further optimize SAM’s segmentation process by incorporating scene geometry, guiding it to generate masks with stronger multi-view consistency, resulting in more accurate features for training 3D language fields.
@@ 2.3 Open-vocabulary Scene Understanding
Scene understanding is a basic task in the field of computer vision. With the rapid advancement of deep learning, numerous methods [40, 4, 5, 37] have made significant progress across various subtasks of scene understanding. However, the limited availability of 3D training data has long posed a challenge for achieving comprehensive 3D scene understanding. Vision foundation models like CLIP [32] have opened new avenues for open-vocabulary scene understanding. Due to the image-aligned nature of CLIP features, subsequent works such as CLIP2Scene [6] and Openscene [29] directly leverage CLIP-based 2D scene understanding models [9, 23] to obtain dense features with CLIP semantics. Approaches like [44, 25, 19, 47, 36] generate dense semantic features by performing multi-level image cropping and feature fusion, while often incorporating pixel-aligned feature supervision, such as DINO [3], to address the blurriness in semantic boundary. Some other researches [21, 14, 16, 30] utilize pre-trained image segmentation model [23, 7] to obtain semantically meaningful object-level patches, enabling more accurate scene understanding.
@@ 3.1 Granularity-aware Segmentation
Observations. The segmentation granularity of SAM is controlled by the prompt point density. In the ramen bowl region of Fig. 3(d)(f), Dense prompt points usually lead to small segmentation regions while sparse prompt points result in large segmentation regions. Another observation is that if the object is far away from the camera, the object is small on the image and the segmentation mask should be small while nearby objects should be large on the image. In this case, the prompt point density is depth-dependent, where we should use a higher prompt point density for distant viewpoints to segment small objects while a lower density for nearby viewpoints to segment the same large objects.
@@
Granularity-aware prompt point density. We compute the prompt point density from the depth maps as follows. As illustrated in Fig. 6, Given an input image, we begin by dividing it into a grid of patches and our target is to determine how many sample points should lie in each patch. Since the rendered depth determines the density, we first render a depth map $D$ from the 3D Gaussian field. Then, we convert the depth map into a set of depth points and reproject these depth points onto all the other visible views. Then, we find the minimum depth values for all the visible views, denoted by $MD$ which maps a pixel to a minimum visible depth value of all visible views. Then, the sample point number for a specific patch $P$ is computed by
$n_{P}=\frac{1}{|P|}\sum_{p\in P}\frac{MD^{2}(p)}{D^{2}(p)}\cdot n,$ (1)
where $n_{P}$ is the number of prompt points in $P$, $|P|$ is the pixel number in the patch, $p\in P$ is a pixel, $MD^{2}(p)$ is the squared minimum depth value of pixel $p$ while $D^{2}(p)$ is the squared depth value of pixel $p$, and $n$ is a predefined prompt point number.
Explanation of Eq. 1. The ratio $\frac{MD^{2}(p)}{D^{2}(p)}\in(0,1]$ controls the prompt point density on a specific view. When this view is the one with the smallest viewing distance, the ratio is 1.0, which means we sample $n$ points on the nearest view. While, for a far-away viewpoint, we use fewer prompt points than the nearest view because the object would be smaller on these far-away views. In this way,  Eq. 1 results in multiview consistent prompting for the SAM model on different viewpoints, which increases the multiview consistency of the segmentation results. Moreover, we find that using random prompt points in the patch produces better results than the even prompt points.
@@ 3.2 Granularity-aware Distillation
Though the above granularity-aware segmentation improves the consistency, the prediction of SAM still exists with randomness and leads to inconsistency. Thus, we further propose a novel granularity-aware distillation to autonomously select the granularity with the best multi-view consistency during optimization. Given the granularity-aware prompt points, SAM returns three levels of segmentation results and on all three levels, we extract CLIP feature maps $f_{s}$, $f_{p}$ and $f_{w}$ corresponding to “sub-part”, “part”, and “whole”. We design a strategy to let the distillation automatically select the most multiview-consistent features.
Decoding granularity factor. Given the 3D Gaussian field associated with trainable feature vectors, we apply the splatting technique to render a feature map for a training viewpoint. We adopt a feature decoder to decode a granularity factor and a predicted CLIP feature from the rendered feature vectors by
@@
where $f_{\text{clip}}$ is the predicted CLIP feature, $\eta\in\mathbb{R}^{3}$ is the predicted granularity factor, $\mathcal{D}$ means the MLP-based decoder shared among all viewpoints, and $f_{\text{render}}$ is the rendered feature.
Weighted distillation loss. We convert the granularity factor into weight by applying the softmax operator between three scores $w=\text{softmax}(\eta)$, which corresponds to the weight of the “sub-part”, “part”, and “whole” respectively. Then, we will train with a weighted L2 loss for distillation
$\ell_{\text{distill}}=\sum_{n\in\{s,p,w\}}w_{n}\|f_{\text{clip}}-f_{n}\|_{2}^{2},$ (3)
where $n$ means the name, $s,p,w$ correspond to “sub-part”, “part”, and “whole” respectively, and $w_{n}\in\mathbb{R}$ means the weight for the current level. We do not directly supervise $w_{n}$ but let the optimization process automatically select the best weight to learn the granularity of the feature.
Entropy regularization. To further encourage the weight to converge to a single scale instead of learning an averaged feature, we adopt an entropy regularization term,
$\ell_{\text{entropy}}=-\sum_{n\in\{s,p,w\}}w_{n}\log w_{n},$ (4)
@@
$\ell_{\text{r-distill}}=\alpha_{r}{\ell_{\text{distill}}},\alpha_{r}=\frac{\sum_{i=1}^{n_{r}}{S}({R}_{i})}{n_{r}\cdot{{S}({R})}},$ (5)
where $\alpha_{r}$ denotes the region-aware factor, with $n_{r}$ as the number of objects in the image and ${S}({R})$ as the area of each region.
@@
$\ell_{\text{cons}}=\sum_{i=1}^{n_{r}}\sum_{p\in R_{i}}\frac{(f_{p}-\overline{f_{R_{i}}})^{2}}{S(R_{i})}$ (6)
where $f_{p}$ and $\overline{f_{R_{i}}}$ represent the $f_{\text{clip}}$ of pixel $p$ and the average $f_{\text{clip}}$ of the $i$-th region. By pulling together features of the same object region across each viewpoint, we ultimately enhance the feature consistency for the 3D Gaussians belonging to the same object.
@@ 4.1 Experimental Protocol
Implementation Details. We utilized the SAM ViT-H [10] and OpenCLIP ViT-B/16 [8] for segmentation and CLIP feature extraction across the multi-view images. All experiments are conducted on a single RTX-4090 GPU. More details on implementation are provided in the supplementary material.
@@ 4.2 Comparison with Baseline Methods
In contrast, our method encourages the distillation of multi-view consistent CLIP features into the Gaussian field, avoiding noise from conflicting multi-view features, which leads to a 4.1% and 3.4% improvement in localization and segmentation accuracy on the LERF dataset. Moreover, the proposed Granularity-aware Distillation can adaptively fuse distill multi-granularity features, ensuring the effectiveness of GAGS in multi-scale object localization. Consequently, on the Mip-NeRF-360 dataset, which includes objects of varying scales, we achieved a more significant improvement over the baseline, with a 15.5% mAcc improvement for localization and a 6.0% increase in mIoU for segmentation.
Runtime analysis. In Tab. 3, we report the runtime of baselines and GAGS. Although GOI [31] achieved the shortest training time, it requires fine-tuning a high-dimensional hyperplane during testing to segment the target, resulting in an inference time nearly 100 times that of GAGS. When comparing with LangSplat [30], GAGS achieves a similar training speed. However, LangSplat requires rendering multi-granularity feature maps to compare and segment the target, while GAGS adaptively fuses multi-granularity features, necessitating only a single feature map rendering and comparison. Thus, GAGS is two times faster than LangSplat in inference.
@@
Table 3: Time evaluation for each step in training and prediction on the scene ”ramen” of the LERF dataset. During training, “Prep.” refers to the preprocessing stage, while “GS” and “Lang.” represent the training stage of RGB and feature Gaussian fields, respectively. During inference, “IE” denotes obtaining the Identity Encoding of the target, while “Rend.” and “Pred.” refer to rendering feature image and predicting target mask.
@@ 4.3 Ablation Studies
In Tab. 4 and Fig. 9, we conduct both qualitative and quantitative ablation evaluations of GAGS on the Mip-NeRF-360 dataset. Model 1 of Tab. 4 is the baseline model, where we use uniform prompt points for SAM segmentation. We then average the CLIP features across different SAM granularities, calculated as $(f_{s}+f_{p}+f_{w})/3$, and distilled the normalized feature into the Gaussian field. Model 2 incorporate the proposed Granularity-aware Distillation (GaD), which boosts the localization and segmentation accuracy by 7.2% and 9.4%, respectively. Further incorporating the proposed Granularity-aware Segmentation (GaS) in Model 3 brings an additional 3.0% mAcc improvement for localization and 2.4% mIoU improvement for segmentation. The results in Fig. 9(a) and (c) also shows that GaS can mitigate noise issues from over-segmentation of nearby objects and low-texture areas, as well as under-segmentation of distant objects.
Moreover, in Models 4-6, we present results by distilling CLIP features using single SAM segmentation granularity. Among them, the part granularity yields the best results, while the subpart granularity performed the worst because the over-segmentation makes CLIP feature extraction unreliable and leads to multi-view conflicts. Model 7, which simply averages the CLIP features from all three SAM granularities, did not yield improvement. In contrast, Model 3, using GaD, can adaptively choose reliable and multi-view consistent features, achieving optimal performance. The comparison between panels (b) and (c) in Fig. 9 also demonstrates that GaD significantly improves segmentation accuracy by distilling multi-view consistent features at the appropriate granularity for each object.
@@
Table 4: Ablation Studies on the Mip-NeRF-360 dataset. “GaS” means the proposed granularity-aware segmentation and “Distill.” denotes the feature distillation strategy. For “Distill.”, in addition to distilling the averaged features across all granularities, we also conducted experiments for each individual granularity feature.

### v2

@@ 1 Introduction
A noticeable challenge of learning such open-vocabulary features for 3D fields lies in the multiview inconsistency of the extracted 2D features. As shown in  Fig. 2, given a specific 3D location, it can be described as “desktop”, “bowl”, and “Ramen” at the same time according to the different granularities. Such semantic uncertainty often results in inconsistent feature extraction across varying viewpoints. Brute-force training a 3D feature field from inconsistent 2D features leads to averaged and degenerated 3D features. LangSplat [33] resolves this problem by learning three different feature fields corresponding to different segmentation granularity of SAM [22]. However, learning three different fields increases both training memory and time cost and requires the query texts to be compared with three different features. FastLGS [18] accelerates open-vocabulary queries by indexing multi-view features of the same object into low-dimensional feature grids. However, it requires additional storage for high-dimensional CLIP features, which scales with the number of views. N2F2 [2] proposes to merge the features of different granularity by selecting the most activated features within a set of manually predefined descriptions. In this case, the performance is strongly related to the quality of these predefined descriptions. If the granularity of query descriptions differs from these predefined descriptions, the performance degenerates severely.
Figure 3: In comparison to vanilla SAM, we associate the prompt point density with distance to the object, which greatly enhances the consistency of multi-view segmentation.
@@
First, we improve the segmentation granularity to improve the consistency. In the CLIP feature extraction, we apply the SAM to segment images into regions for dense feature extraction, but its multi-solution nature causes viewpoint inconsistency. As shown in  Fig. 3, we improve the consistency of SAM segmentation results among multiview images by introducing a granularity-aware prompting strategy. SAM relies on the prompt points to determine the segmentation granularity. Thus, we associate the prompt point density with the camera distance to the subject by utilizing the reconstructed 3D GS. Distant views receive dense prompts while nearby views use sparse prompts. We found that this adaptive prompting effectively results in consistent segmentation across multiview images and further improves the CLIP feature consistency.
@@ 2.2 SAM
As one of the most impressive foundational vision models, Meta’s SAM [22] has demonstrated exceptional zero-shot 2D segmentation capabilities. SAM supports flexible prompts, allowing it to generate multi-level masks for target objects based on inputs such as points, bounding boxes, and masks. Additionally, SAM offers the capability to automatically segment the entire image into multi-granularity masks, including whole, subpart and part. Numerous methods have since emerged to extend SAM’s capabilities to 3D space. Anything-3D [37] elevates SAM’s segmentation to 3D, while Part123 [27] uses SAM’s segmentation masks to reconstruct 3D models with high-quality segmented parts. Feature 3DGS [51] and Gaussian Grouping [46] integrate SAM’s features and object segmentation results at pre-selected single-granularity into 3D Gaussians, achieving high-quality segmentation of novel views and 3D scenes. Semantic Gaussian [13] utilizes various prompts to obtain instance-level segmentations for feature extraction.
Considering that SAM can produce multi-granularity masks focusing on objects at different scales, some methods [21, 48, 26] integrates these masks into scene representations jointly or separately. However, the above methods, whether single-granularity or multi-granularity, do not address the issue of multi-view mask conflicts. In contrast, we enhance the consistency and reliability of multi-view masks in SAM segmentation and mask integration process.
@@ 2.3 Open-vocabulary Scene Understanding
Scene understanding is a basic task in the field of computer vision. With the rapid advancement of deep learning, numerous methods [44, 4, 5, 40] have made significant progress across various subtasks of scene understanding. However, the limited availability of 3D training data has long posed a challenge for achieving comprehensive 3D scene understanding. Some methods [43, 46] attempt to align the outcomes of 2D grounding models [36] with 3D category-agnostic grouping approaches, thereby indirectly achieving open-vocabulary understanding. Vision foundation models like CLIP [35] have opened new avenues for open-vocabulary scene understanding. Due to the image-aligned nature of CLIP features, subsequent works such as CLIP2Scene [6] and Openscene [32] directly leverage CLIP-based 2D scene understanding models [9, 25] to obtain dense features with CLIP semantics. Approaches like [49, 28, 20, 52, 39] generate dense semantic features by performing multi-level image cropping and feature fusion, while often incorporating pixel-aligned feature supervision, such as DINO [3], to address the blurriness in semantic boundary. Some other researches [23, 14, 16, 33] utilize pre-trained image segmentation model [25, 7] to obtain semantically meaningful object-level patches, enabling more accurate scene understanding.
@@ 3.1 Granularity-aware Segmentation
Observations. The segmentation granularity of SAM is controlled by the prompt point density. In the ramen bowl region of Fig. 3(d)(f), dense prompt points usually lead to small segmentation regions while sparse prompt points result in large regions. Another observation is that objects far from the camera appear smaller in the image, requiring smaller segmentation masks, while nearby objects appear larger and require larger masks. In this case, prompt point density becomes depth-dependent: higher density is needed for distant viewpoints to segment smaller objects, while lower density suffices for nearby viewpoints to segment larger objects.
@@
Granularity-aware prompt point density. We compute the prompt point density from the depth maps as follows. As illustrated in Fig. 6, Given an input image, we begin by dividing it into a grid of patches and our target is to determine how many sample points should lie in each patch. Since the rendered depth determines the density, we first render a depth map $D$ from the 3D Gaussian field. Then, we back-project each depth map pixel to the 3D Gaussians and find the minimum depth value of each Gaussian after occlusion determination, denoted by $MD$, which maps a pixel to the minimum visible depth across all views. Then, the number of prompt points $n_{P}$ for a specific patch $P$ is computed by
$n_{P}=\frac{1}{|P|}\sum_{p\in P}\frac{D^{2}(p)}{MD^{2}(p)}\cdot n,$ (1)
where $|P|$ is the pixel number in the patch, where $|P|$ is the pixel number in the patch, $p\in P$ is a pixel, $MD^{2}(p)$ is the squared minimum depth value of pixel $p$ while $D^{2}(p)$ is the squared depth value of pixel $p$, and $n$ is a predefined prompt point number.
Explanation of Eq. 1. The ratio $\frac{D^{2}(p)}{MD^{2}(p)}$ controls the prompt point density on a specific view. When this view is the one with the smallest viewing distance, the ratio is 1.0, which means we sample $n$ points on the nearest view. While, for a far-away viewpoint, we use a larger prompt point density because the object would be smaller on these views, which can be seen in Fig. 3(e)(f). In this way,  Eq. 1 enables granularity-invariant prompting for the SAM model when segmenting the same object from different viewpoints, which increases the multi-view consistency of the segmentation results.
Moreover, after determining the number of prompt points per patch, the density distribution of visible gaussians is utilized to guide local prompt point sampling. Please refer to the supplementary material for further details.
@@ 3.2 Granularity-aware Distillation
While the above granularity-aware segmentation improves the consistency, SAM’s prediction still exists with randomness and leads to inconsistency. Thus, we further propose a novel granularity-aware distillation to autonomously select the granularity with the best multi-view consistency during optimization. Given the granularity-aware prompt points, SAM returns three levels of segmentation results $m_{s}$, $m_{p}$ and $m_{w}$ corresponding to “sub-part”, “part”, and “whole”. We then extract CLIP feature maps $f_{s}$, $f_{p}$ and $f_{w}$ on all three levels and design a strategy to let the distillation automatically select the most multiview-consistent features.
Decoding granularity factor. Given the 3D Gaussian field associated with trainable feature vectors, we apply the splatting technique to render a feature map $f_{\text{render}}$ for a training viewpoint. We adopt a feature decoder to decode a granularity factor and a predicted CLIP feature from the rendered feature vectors by
@@
where $f_{\text{clip}}$ is the predicted CLIP feature, $\eta\in\mathbb{R}^{3}$ is the predicted granularity factor, and $\mathcal{D}$ means the MLP-based decoder shared among all viewpoints.
Weighted distillation loss. We convert the granularity factor into weight by applying the softmax operator between three scores $\alpha=\text{softmax}(\eta)$. Then, we will train with a weighted L2 loss for distillation
$\ell_{\text{distill}}=\sum_{n\in\{s,p,w\}}\alpha_{n}\|f_{\text{clip}}-f_{n}\|_{2}^{2},$ (3)
where $n$ means the name, $s,p,w$ correspond to “sub-part”, “part”, and “whole” respectively, and $\alpha_{n}\in[0,1]$ means the weight for the current level. We do not directly supervise $\alpha_{n}$ but let the optimization process automatically select the best weight to learn the granularity of the feature.
Entropy regularization. To further encourage the weight to converge to a single scale instead of learning an averaged feature, we adopt an entropy regularization term
$\ell_{\text{entropy}}=-\sum_{n\in\{s,p,w\}}\alpha_{n}\log\alpha_{n},$ (4)
@@
$\ell_{\text{r-distill}}=\beta_{r}{\ell_{\text{distill}}},\beta_{r}=\frac{\sum_{i=1}^{n_{r}}{S}({R}_{i})}{n_{r}\cdot{{S}({R})}},$ (5)
where $\beta_{r}$ denotes the region-aware factor, with $n_{r}$ as the number of objects in the image and ${S}({R})$ as the area of each region. However, region-aware strategy is non-trivial. The core challenge here is constructing a mask that can simultaneously segment both large and small objects for balanced supervision, where SAM masks at any single granularity fails to achieve this as the whole scale mask $m_{w}$ focuses on large objects, while the part $m_{p}$ and subpart $m_{s}$ masks emphasize small objects. Therefore, we propose integrating SAM masks at multiple granularities to adaptively select the appropriate scale for objects of different sizes. Specifically, we construct $m_{f}$, where each pixel selects the SAM granularity with the highest weight, i.e., $m_{f}=m_{\underset{i\in\{s,p,w\}}{\arg\max}(\alpha_{i})}$. We then cluster pixels in $m_{f}$ to objects for computing $S(R)$. Further details are provided in the supplementary material.
@@
$\ell_{\text{cons}}=\sum_{i=1}^{n_{r}}\sum_{p\in R_{i}}\frac{(f_{clip}^{p}-\overline{f_{R_{i}}})^{2}}{S(R_{i})}$ (6)
where $f_{clip}^{p}$ and $\overline{f_{R_{i}}}$ represent the $f_{\text{clip}}$ of pixel $p$ and the average $f_{\text{clip}}$ of the $i$-th region. By pulling together features of the same object region across each viewpoint, we ultimately enhance the feature consistency for the 3D Gaussians belonging to the same object.
@@ 4.1 Experimental Protocol
Implementation Details. We utilized the SAM ViT-H [10] and OpenCLIP ViT-B/16 [8] for segmentation and CLIP feature extraction. All experiments are conducted on a single RTX-4090 GPU. More details on implementation are provided in the supplementary material.
@@ 4.2 Comparison with Baseline Methods
In contrast, our method encourages the distillation of multi-view consistent CLIP features into the Gaussian field, avoiding noise from conflicting multi-view features, which leads to a 4.1% improvement in localization mAcc and 3.4% in segmentation mIoU on the LERF dataset. Moreover, the proposed Granularity-aware Distillation can adaptively fuses multi-granularity features, thus ensuring effective multi-scale object localization. Consequently, on the Mip-NeRF-360 dataset, which includes objects of varying scales, we achieved a more significant 15.5% mAcc and 6.0% mIoU improvement over the baseline.
@@
Table 3: Time evaluation for each step in training and prediction on the scene “ramen” of the LERF dataset. During training, “Prep.” refers to the preprocessing stage, while “GS” and “Lang.” represent the training stage of RGB and feature Gaussian fields, respectively. During inference, “IE” and “OSH” denotes obtaining the target’s Identity Encoding and the optimization process of the Optimizable Semantic-space Hyperplane, while “Rend.” and “Pred.” refer to rendering features and predicting target mask.
Runtime analysis. We report the runtime of baselines and GAGS in Tab. 3. Although GOI [34] achieved the shortest training time, it requires fine-tuning a high-dimensional hyperplane during testing to segment the target, resulting in an inference time nearly 100 times that of GAGS. When comparing with LangSplat [33], GAGS achieves a similar training speed. However, LangSplat requires rendering multi-granularity feature maps to compare and segment the target, while GAGS adaptively fuses multi-granularity features, necessitating only a single feature map rendering and comparison. Thus, GAGS is two times faster than LangSplat in inference.
Figure 10: Visualization of prompt points and segmentation results. Our method achieves superior segmentation consistency with only ~20% of the prompt point count used in the vanilla SAM.
Figure 11: Visualization of learned granularity weights $\alpha$. The RGB channels correspond to $\alpha_{s}$, $\alpha_{p}$, and $\alpha_{w}$, respectively.
@@ 4.3 Ablation Studies
In Tab. 4 and Fig. 9, we conduct both qualitative and quantitative ablation evaluations of GAGS on the Mip-NeRF-360 dataset. Model 1 of Tab. 4 is the baseline model, where we use uniform prompt points for SAM segmentation. We then average the CLIP features, calculated as $(f_{s}+f_{p}+f_{w})/3$, and distill the normalized feature into Gaussian field. Model 2 incorporate the proposed Granularity-aware Distillation (GaD), which boosts the localization mAcc and segmentation mIoU by 9.4% and 7.2%, respectively. Further incorporating the proposed Granularity-aware Segmentation (GaS) in Model 3 brings an additional 2.4% mAcc and 3.5% mIoU improvement. As shown in Fig. 10, GaS achieves better multiview segmentation consistency by prompting SAM with fewer but proper seeds. The results in Fig. 9(a) and (c) also shows that GaS can mitigate noise issues from over-segmentation of nearby objects and low-texture areas, as well as under-segmentation of distant objects.
@@
Table 4: Ablation Studies on the Mip-NeRF-360 dataset. “Distill.” denotes the feature distillation strategy. For “Distill.”, in addition to distilling the averaged features across all granularities, we also conducted experiments for each individual granularity feature.
Moreover, in Models 4-6, we distill CLIP features using single SAM segmentation granularity. Among them, the part granularity yields the best results, while the subpart granularity performed the worst because the over-segmentation makes CLIP feature extraction unreliable and leads to multi-view conflicts. Model 7, which simply averages the CLIP features from all three SAM granularities, shows no improvement. In contrast, Model 3 with GaD can adaptively select reliable and multi-view consistent features. As shown in Fig. 11, GaD facilitates the model in selecting features of corresponding granularities for objects of varying scales, ensuring the consistency and robustness of object feature learning. Finally, model 3 achieves optimal performance. The comparison between panels (b) and (c) in Fig. 9 also demonstrates that GaD significantly improves segmentation accuracy by distilling multi-view consistent features at the appropriate granularity for each object.